OpenStudy (anonymous):

What is (5x)(14x^3)?

OpenStudy (btaylor):

it is (5 x 14)(x)(x^3)=70x^4

OpenStudy (anonymous):

TY.

OpenStudy (jackellyn):

5*14= 70 And x(x^3) adds the exponents so 1+3=4 So 70x^4